OpenStudy (anonymous):

Simplify. −3(9x − 4y)(3x − 7y) do i need to multiply the -3 to both or just the first?

OpenStudy (kira_yamato):

Just the first...

OpenStudy (anonymous):

ok then after that add the like terms?

OpenStudy (anonymous):

use the (a-b)^2 =a^2-b^2

OpenStudy (zzr0ck3r):

-3(27x^2-7*9*x*y-4*3*x*y+28*y^2)= -3*27*x^2+3*7*9*x*y+3*3*4*x*y-3*28*y^2

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

dont do what rohangrr said (a-b)^2 = (a-b)*(a+b)

OpenStudy (zzr0ck3r):

foil the binomials then factor the -3 through or factor the -3 into one of the binomials and then foil

OpenStudy (anonymous):

ok